# Monte Carlo Tree Diffusion for System 2 Planning

### v6

@@ 1 Introduction
In contrast, Monte Carlo Tree Search (MCTS) (Coulom, 2007), a widely adopted planning method, demonstrates robust inference-time scalability. By leveraging iterative simulations, MCTS refines decisions and adapts based on exploratory feedback, making it highly effective in improving planning accuracy as more computation is allocated. This capability has established MCTS as a cornerstone in many System 2 reasoning tasks, such as mathematical problem-solving (Guan et al., 2025; Zhang et al., 2024a) and program synthesis (Brandfonbrener et al., 2024). However, unlike diffusion-based planners, traditional MCTS relies on a forward model for tree rollouts, inheriting its limitations including losing global consistency. In addition to being restricted to discrete action spaces, the resulting search tree can grow excessively large in both depth and width. This leads to significant computational demands, particularly in scenarios involving long horizons and large action spaces.
@@ 2.1 Diffuser: Diffusion Models for Planning
Accordingly, at each denoising step, gradient information from nudges the model toward trajectories that appear both feasible (as learned from the offline data) and promising with respect to returns, in a manner akin to classifier guidance in image diffusion (Dhariwal & Nichol, 2021).
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree search with stochastic simulations to effectively balance exploration and exploitation (Coulom, 2007). Typically, MCTS proceeds in four stages: selection, expansion, simulation, and backpropagation. During selection, the algorithm starts at a root node corresponding to the current state (or partial plan) and descends the tree according to a policy such as Upper Confidence Bounds for Trees (UCT) (Kocsis & Szepesvári, 2006), aiming to choose the most promising child at each step or explore unexperienced states for balanced. Once a leaf or expandable node is reached, expansion adds new child nodes representing previously unexplored actions or plans.
@@ 3.2 Guidance Levels as Meta-Actions
For simplicity, consider two guidance levels: guide and no_guide. In MCTD, we observe that sampling from the prior distribution , i.e., using the standard diffusion sampler represents exploratory behavior as it does not attempt to achieve any goal but only imitates the prior behavior contained in the offline data. Thus, we correspond it to the no_guide meta-action. Conversely, sampling from a goal-directed distribution , such as through classifier-guided diffusion (Dhariwal & Nichol, 2021), represents exploitative behavior, and thus meta-action, guide, is assigned. This steers the sampling process toward achieving a specific goal defined by a reward function .
@@ 4 Related Works
MCTS (Coulom, 2007) has achieved impressive results across various decision-making problems, particularly when combined with learned policies or value networks (Silver et al., 2016; Schrittwieser et al., 2020). It has also been applied to System 2 reasoning in Large Language Models (LLMs) to enhance structured problem-solving (Xiang et al., 2025; Yao et al., 2024; Zhang et al., 2024b). However, to the best of our knowledge, this work is the first to integrate tree search with diffusion models for full trajectory generation, bridging structured search with generative planning.
@@ 5.4 Visual Pointmaze
We pre-encode observations with a Variational Autoencoder (VAE) (Kingma & Welling, 2014) to generate latent states suitable for diffusion-based planning, . Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, . For guidance, we employ a position estimator trained on ground-truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. Further details on this task are provided in Appendix A.7.
@@ A.7.1 Setup
Since diffusion-based planners typically operate on state-action representations, we first learn an image encoder using a Variational Autoencoder (VAE) (Kingma & Welling, 2014). The resulting latent states serve as the input to our diffusion model. Notably, these latent encodings do not contain explicit positional information; consequently, the planner must infer the underlying geometry to navigate the maze indirectly. We design an inverse dynamics model—a compact Multi-Layer Perceptron (MLP)—to map consecutive latent observations to actions. This model replaces the heuristic controller from the original Diffuser architecture, which is not directly applicable in this partially observable setting.
@@ A.7.5 Variational Autoencoder
We pretrain a VAE (Kingma & Welling, 2014) to encode RGB observations into an -dimensional latent space. This compresses the high-dimensional visual inputs into a structured representation suitable for downstream planning. The VAE models the latent variable as a Gaussian distribution with mean and standard deviation , using the reparameterization trick:
@@ D.1 UCT Hyperparameter Ablation Study
In addition to the meta-action ablation study (Section 5.6.2), we empirically analyze the exploration-exploitation balance by varying the UCT (Kocsis & Szepesvári, 2006) hyperparameter in the formula: , where is the node’s estimated value, and and are the visit counts of the parent node and node itself, respectively.

### v7

@@ 1 Introduction
In contrast, Monte Carlo Tree Search (MCTS) (Coulom, 2007), a widely adopted planning method, demonstrates robust inference-time scalability. By leveraging iterative simulations, MCTS refines decisions and adapts based on exploratory feedback, making it highly effective in improving planning accuracy as more computation is allocated. This capability has established MCTS as a cornerstone in many System 2 reasoning tasks, such as mathematical problem-solving (Guan et al., 2025; Zhang et al., 2024b) and program synthesis (Brandfonbrener et al., 2024). However, unlike diffusion-based planners, traditional MCTS relies on a forward model for tree rollouts, inheriting its limitations including losing global consistency. In addition to being restricted to discrete action spaces, the resulting search tree can grow excessively large in both depth and width. This leads to significant computational demands, particularly in scenarios involving long horizons and large action spaces.
@@ 2.1 Diffuser: Diffusion Models for Planning
Accordingly, at each denoising step, gradient information from nudges the model toward trajectories that appear both feasible (as learned from the offline data) and promising with respect to returns, in a manner akin to classifier guidance in image diffusion (Dhariwal and Nichol, 2021).
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree search with stochastic simulations to effectively balance exploration and exploitation (Coulom, 2007). Typically, MCTS proceeds in four stages: selection, expansion, simulation, and backpropagation. During selection, the algorithm starts at a root node corresponding to the current state (or partial plan) and descends the tree according to a policy such as Upper Confidence Bounds for Trees (UCT) (Kocsis and Szepesvári, 2006), aiming to choose the most promising child at each step or explore unexperienced states for balanced. Once a leaf or expandable node is reached, expansion adds new child nodes representing previously unexplored actions or plans.
@@ 3.2 Guidance Levels as Meta-Actions
For simplicity, consider two guidance levels: guide and no_guide. In MCTD, we observe that sampling from the prior distribution , i.e., using the standard diffusion sampler represents exploratory behavior as it does not attempt to achieve any goal but only imitates the prior behavior contained in the offline data. Thus, we correspond it to the no_guide meta-action. Conversely, sampling from a goal-directed distribution , such as through classifier-guided diffusion (Dhariwal and Nichol, 2021), represents exploitative behavior, and thus meta-action, guide, is assigned. This steers the sampling process toward achieving a specific goal defined by a reward function .
@@ 4 Related Works
MCTS (Coulom, 2007) has achieved impressive results across various decision-making problems, particularly when combined with learned policies or value networks (Silver et al., 2016; Schrittwieser et al., 2020). It has also been applied to System 2 reasoning in Large Language Models (LLMs) to enhance structured problem-solving (Xiang et al., 2025; Yao et al., 2024; Zhang et al., 2024a). However, to the best of our knowledge, this work is the first to integrate tree search with diffusion models for full trajectory generation, bridging structured search with generative planning.
@@ 5.4 Visual Pointmaze
We pre-encode observations with a Variational Autoencoder (VAE) (Kingma and Welling, 2014) to generate latent states suitable for diffusion-based planning, . Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, . For guidance, we employ a position estimator trained on ground-truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. Further details on this task are provided in Appendix A.7.
@@ A.7.1 Setup
Since diffusion-based planners typically operate on state-action representations, we first learn an image encoder using a Variational Autoencoder (VAE) (Kingma and Welling, 2014). The resulting latent states serve as the input to our diffusion model. Notably, these latent encodings do not contain explicit positional information; consequently, the planner must infer the underlying geometry to navigate the maze indirectly. We design an inverse dynamics model—a compact Multi-Layer Perceptron (MLP)—to map consecutive latent observations to actions. This model replaces the heuristic controller from the original Diffuser architecture, which is not directly applicable in this partially observable setting.
@@ A.7.5 Variational Autoencoder
We pretrain a VAE (Kingma and Welling, 2014) to encode RGB observations into an -dimensional latent space. This compresses the high-dimensional visual inputs into a structured representation suitable for downstream planning. The VAE models the latent variable as a Gaussian distribution with mean and standard deviation , using the reparameterization trick:
@@ Appendix B MCTD Architecture Implementation
Additionally, when expanding from the root node to multiple child nodes, each child node begins denoising from a different initial noise sampled independently from a Gaussian distribution. By allowing each child to start from a distinct initial noise, we leverage the tree search structure more effectively. Rather than searching only within the trajectory distribution conditioned on a single initial noise, MCTD explores across multiple initial noises, enabling a broader search over diverse trajectory modes.
@@ D.1 UCT Hyperparameter Ablation Study
In addition to the meta-action ablation study (Section 5.6.2), we empirically analyze the exploration-exploitation balance by varying the UCT (Kocsis and Szepesvári, 2006) hyperparameter in the formula: , where is the node’s estimated value, and and are the visit counts of the parent node and node itself, respectively.